OpenStudy (vampirediaries):

A cannery processed 1,445 pounds of strawberries in 4.5 hours. The cannery processed 2,320 of strawberries in 7 hours. Write a linear equation to model the relationship between the weight of strawberries S and time T. How many pounds of strawberries can be processed in 11 hours?

OpenStudy (vampirediaries):

@cj49, @aum, @kropot72 can you help me, please?

OpenStudy (cj49):

2320 7 x 11 11x2320/7=?

OpenStudy (vampirediaries):

I got 3645.71. What's the next step?

OpenStudy (cj49):

thats the ans in 11 hrs it wud process 3645.71 pounds

OpenStudy (vampirediaries):

Oohh ok. Wow that was easier than I expected. Thank you for help

OpenStudy (cj49):

np c:

OpenStudy (vampirediaries):

Wait... How would the equation look like?

OpenStudy (cj49):

i guess its s/t wud give u the amt of strawberries processed in an hr.2320/7=331.42 331.42 pounds is processed per hr. Now foh 11 hrs =11*331.42

OpenStudy (vampirediaries):

Okay

OpenStudy (aum):

"Write a linear equation to model the relationship between the weight of strawberries S and time T". Linear equation means you can assume that if you plot the weight of the strawberries S against time T, the plot will be a straight line. You are given two points on the line from which you can find the equation of the line. The two points are: (4.5, 1445) and (7, 2320). Slope = (2320-1445) / (7-4.5) = 875 / 2.5 = 350. S = mT + b S = 350T + b put one of the points and find b 1445 = 350*4.5 + b = 1575 + b b = 1445-1575 = -130 S = 350T - 130 when T = 11 hours S = 350*11 - 130 = 3720 pounds.